Session: Mar 2008 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Tao An at Fri Feb 29 04:36:21 2008) Minor failures : RECO
1.the warning "recording throttled can't keep up " for everyscan. 2.the error "scan_check" failed for some scans.

Urumqi (Hua Zhang at Fri Feb 29 06:59:24 2008) Success :

Torun (Kaz Borkowski at Fri Feb 29 07:00:20 2008) Success :
No known problems.

Medicina (Fiocchi Franco at Fri Feb 29 07:03:28 2008) Success :

Hartebeesthoek (Jonathan Quick at Fri Feb 29 07:33:27 2008) Failure : RECE HIGH
Cable break on antenna caused expander to stop resulting in cryogenic package warmup. Unfortunately the LNAs oscillate at room temperature so calibration work is impossible. We ran the experiment anyway, but don't expect the data to be useable.

Onsala (Michael Lindqvist at Fri Feb 29 12:25:18 2008) Minor failures : TSYS
Could not read zero-offset (tpzero) for VC02. Thus, could not get Tsys-values for VC02. RFI.

Effelsberg (Uwe Bach at Mon Mar 3 12:12:27 2008) Success :

Jodrell_Bank (Alastair Gunn at Mon Mar 3 15:45:43 2008) Success :
The Lovell Telescope was used (Jb1). No known problems.

Noto (P. Cassaro at Tue Mar 4 17:07:23 2008) Minor failures : TSYS
The lower side band of the bbc #8 was in overflow for the whole experiment.

Westerbork (Geert Kuper at Mon Mar 17 15:45:48 2008) Success :
Wb used in phased array mode.